# FlightScope: A Deep Comprehensive Review of Aircraft Detection Algorithms in Satellite Imagery

### v3

@@ 4.1 You Only Look Once
The You Only Look Once (YOLO) framework [59] has emerged as a popular deep learning-based object detection algorithm that revolutionized real-time object detection tasks. It presents a unified approach to object detection by formulating it as a regression problem, enabling the model to predict bounding boxes and class probabilities of multiple objects in a single pass through the network.
@@
In the early stages, YOLOv2 [63] introduced batch normalization and high-resolution classifiers to improve detection performance. YOLOv3 [33] further refined the architecture by incorporating skip connections and multi-scale prediction, enhancing both accuracy and localization capabilities. The introduction of CSPDarknet, SPP, PAN, and Mish activation in YOLOv4 [35] led to significant improvements in the network architecture. YOLOv5 [64], building upon the scaled-YOLOv4 [65], introduced anchor-free object detection and a new architecture. This version expanded the options available by providing models of varying sizes, allowing users to balance speed and accuracy according to their specific requirements (Fig. 2).
The latest known versions, YOLOv8 [22] and YOLO-NAS [66], brought forth the concept of neural architecture search (NAS) [67] to automatically design network architectures and achieve state-of-the-art performance in object detection tasks. The continuous evolution of the YOLO framework highlights the trade-offs between speed and accuracy, necessitating consideration of the specific application requirements when selecting an appropriate YOLO model.
4.2 Single Shot Detection
@@
The Single Shot Detector (SSD) [68] is an efficient and accurate object detection algorithm that introduces a unified framework for single-pass detection. In theory, SSD addresses the challenge of detecting objects at various scales and aspect ratios by utilizing predefined anchor boxes [69]. The SSD architecture consists of three main components: a base network, convolutional feature maps, and convolutional predictors (Fig. 3). The base network, usually a pre-trained CNN [70], acts as a feature extractor, generating high-level feature maps with different spatial resolutions. These feature maps are then processed by convolutional predictors, which predict object presence and locations for the anchor boxes. Each predictor corresponds to a specific feature map and produces class scores and bounding box offsets. The anchor boxes, distributed densely across the feature map, serve as reference boxes for object detection. To capture objects at multiple scales, SSD employs feature maps from different stages of the base network. Higher-resolution feature maps are effective at detecting small objects, while lower-resolution ones are suitable for larger objects. The predictions from each feature map are combined to generate final class predictions and refined bounding box coordinates. During training, SSD utilizes a multi-task loss function [71] that optimizes the model by considering both localization loss (Smooth L1 loss) and classification loss (softmax loss) [72]. The localization loss penalizes the discrepancy between predicted and ground truth bounding box coordinates, while the classification loss encourages accurate class predictions. The loss is computed for positive and negative samples, including hard negatives based on confidence scores.
@@
While SSD offers a good balance between accuracy and efficiency with its varying feature map resolutions and fixed anchor boxes, it does have limitations [73]. One limitation is the dependence on predefined anchor boxes, which may not adequately cover objects with extreme aspect ratios or unconventional shapes. Scaling the anchor boxes can help address this, but it introduces additional computational overhead. Another limitation is the challenge of handling tiny objects [74]. Since SSD relies on a limited number of feature maps, it may struggle to accurately detect small objects due to the loss of fine-grained details at higher resolutions. This can result in reduced localization accuracy and increased false negatives for small or densely packed objects. Furthermore, SSD’s fixed anchor boxes limit its ability to handle objects at arbitrary scales and aspect ratios.
4.3 Region-based CNNs
@@
The introduction of the Region-based Convolutional Neural Network (R-CNN) [75] marked a significant milestone in the development of object detection techniques, showcasing the substantial improvements that convolutional neural networks (CNNs) can bring to detection performance. R-CNN introduced the concept of utilizing CNNs [76] in combination with a class-agnostic region proposal [77] module to transform object detection into a classification and localization problem. The detection process starts with a mean-subtracted input image, which is fed through the region proposal (RPN) module. This module employs techniques such as Selective Search [78] to identify regions within the image that have a higher likelihood of containing objects. Approximately 2000 object candidates are generated based on this region proposal step. These candidates are then warped and passed through a CNN network, such as the widely used ImageNet [79], to extract a 4096-dimensional feature vector for each proposal. The feature vectors obtained from the CNN are then inputted into class-specific Support Vector Machines (SVMs) [80], which have been trained beforehand. The SVMs generate confidence scores for each candidate region, aiding in the classification process. To refine the results, non-maximum suppression (NMS) is applied based on the Intersection over Union (IoU) and class information. Once the class has been identified, a trained bounding-box regressor is employed to predict the precise bounding box coordinates, including the center coordinates, width, and height of the object (Fig. 4).
@@
However, despite its groundbreaking contributions, R-CNN has several limitations [81, 82]. The training process of R-CNN is complex and multistage. It involves pretraining the CNN with a large classification dataset, followed by fine-tuning on domain-specific images that undergo mean subtraction and warping to align with the proposals. The CNN classification layer is replaced with a randomly initialized N+1-way classifier, where N represents the number of classes, and stochastic gradient descent (SGD) is utilized for optimization. Additionally, separate SVMs and bounding box regressors need to be trained for each class, adding to the computational complexity.
@@
Although R-CNN is capable of performing highly accurate object detection research, it suffered from slow inference times, taking approximately 47 seconds per image, and was resource intensive in terms of both time and space [75]. Training R-CNN models on small datasets took days to complete, even with shared computations. These limitations sparked the need for further advancements in object detection algorithms that could address these challenges and improve overall efficiency.
4.4 RetinaNet Framework
RetinaNet is a one-stage object detection model that was introduced by Tsung-Yi Lin et. al [26] designed to address the extreme foreground-background class imbalance encountered during training of dense detectors. In object detection, the goal is to detect objects of interest in an image and localize them by drawing bounding boxes around them. However, the vast majority of the image is typically background, and there are usually many more negative examples (background) than positive examples (objects of interest). This class imbalance can make it difficult for the detector to learn to distinguish between objects and background, and can lead to poor performance. RetinaNet is a single, unified network that is composed of a backbone network and two task-specific subnetworks (Fig. 5). The backbone is responsible for computing a convolutional feature map over an entire input image and is an off-the-shelf convolutional network. The first subnetwork is a dense prediction subnet that produces a fixed number of object detections of different scales and aspect ratios at each spatial location in a feature map. The second subnetwork is a set of class-specific subnets that further refine the predictions of the first subnet.
Figure 5: One stage RetinaNet architecture. Reproduced from [17]. (a) ResNet [83] backbone. (b) generation of multiscale convolutional pyramid. This is attached to two subnetworks: (c) anchor box classification and (d) anchor box regression to ground truth bounding box.
This model uses a novel focal loss function that down-weights the contribution of easy examples during training to focus on hard examples and prevent the vast number of easy background examples from overwhelming the detector. The focal loss is a dynamically scaled cross-entropy loss, where the scaling factor decays to zero as confidence in the correct class increases. This loss function is designed to address the mechanisms used by other detectors to address class imbalance, such as biased minibatch sampling and object proposal mechanisms, in a one-stage detection system directly via the loss function.
The obtained results showed a good performance compared to previous one-stage and two-stage detectors, including the best reported Faster R-CNN system, on the COCO dataset. It achieves state-of-the-art performance on both the COCO detection tasks, with a better COCO test-dev average precision while running at 5 fps. This is a significant improvement over previous state-of-the-art techniques for training one-stage detectors, such as training with the sampling heuristics or hard example mining. RetinaNet is also designed to be efficient and scalable. It uses an efficient in-network feature pyramid that allows it to detect objects at multiple scales and resolutions, and it uses anchor boxes to improve localization accuracy. The anchor boxes are pre-defined boxes of different scales and aspect ratios that are placed at each spatial location in the feature map. The dense prediction subnet predicts the offsets and scales of the anchor boxes to generate object detections.
4.5 CenterNet Framework
CenterNet [18] is another real-time object detection algorithm designed to operate in real-time, with an average inference time of 270 ms using a 52-layer hourglass backbone and 340 ms using a 104-layer hourglass backbone per image according to the author (Fig. 6). CenterNet is inspired from the architecture of CornerNet [25] which is based on a one-stage keypoint-based detector, while introducing several novel components and strategies to enhance its effectiveness.
Figure 6: One stage CenterNet architecture. Reproduced from [18].
Unlike CornerNet, which detects object bounding boxes using pairs of keypoints, CenterNet introduces the concept of detecting each object as a triplet of keypoints. This innovation allows CenterNet to capture more comprehensive information about the objects, leading to improved detection performance while being fast for live inferencing. CenterNet incorporates two customized modules named cascade corner pooling and center pooling. These modules play crucial roles in enriching information collected by both top-left and bottom-right corners and providing more recognizable information at the central regions, respectively. These modules contribute to the improved performance of CenterNet by enhancing the detection of object keypoints and bounding boxes. In their benchmark study [18], authors stated that this proposed architecture achieves significant improvements over existing one-stage detectors typically an average precision of 47.0% on the MS-COCO dataset, outperforming the previously proposed one-stage detectors by at least 4.9%. Additionally, it demonstrates comparable performance to the two-stage detectors while maintaining a faster inference speed thanks to the effective reduction of incorrect bounding boxes, particularly for small objects. It achieves notable improvements in the detection of small objects, with an average precision (AP) improvements of 5.5% to 8.1% for different backbone configurations. This reduction in incorrect bounding boxes is attributed to the effectiveness of CenterNet in modeling center information using center keypoints.
4.6 End-to-End Transformer
@@
The advent of transformer architectures has revolutionized the field of artificial intelligence by introducing a novel approach to processing sequential data. Unlike older neural network architectures, transformers rely on self-attention mechanisms [44] to capture dependencies between input tokens, enabling them to effectively model long-range dependencies and capture complex patterns in the data. This ability to process sequences in parallel, rather than sequentially, has significantly improved the efficiency and effectiveness specially of natural language processing (NLP) tasks [84]. The success of transformers can be attributed to their ability to capture global context and relationships within the input data, making them particularly well-suited for tasks that require understanding of complex interdependencies.
@@
Later-on, researchers have translated the use of transformers in NLP to computer vision including object detection operation. One pioneering algorithm is DETR (DEtection TRansformer) [52] belonging to the category of end-to-end object detection systems based on transformers and bipartite matching loss for direct set prediction. Unlike traditional object detection methods, DETR does not fall into the conventional one-stage or two-stages detector categories. Instead, it introduces a novel approach by predicting all objects at once using a bipartite matching loss function, which uniquely assigns a prediction to a ground truth object and is invariant to a permutation of predicted objects. This unique approach simplifies the detection pipeline by eliminating the need for hand-designed components such as spatial anchors or non-maximal suppression [85], making it optimal for both accuracy and real-time processing. DETR achieves this by leveraging transformers with parallel decoding, as opposed to autoregressive decoding with recurrent neural networks, which was the focus of previous work.
@@
The architecture of DETR is simple yet highly effective. It comprises three main components: a CNN backbone for feature extraction [86], an encoder-decoder transformer [87] for modeling relationships between feature representations of different detections, and a simple feed-forward network for making the final detection predictions (Fig. 7). The process begins with the extraction of a lower-resolution activation map from the input image using a conventional CNN backbone. This activation map is then passed through the transformer encoder, along with spatial positional encodings that are added to queries and keys at every multi-head self-attention layer. The decoder receives queries, output positional encodings (object queries), and encoder memory, and produces the final set of predicted class labels and bounding boxes through multiple multi-head self-attention and decoder-encoder attention (FFNs). The simplicity and modularity of the DETR architecture make it easily implementable in any deep learning framework that provides a common CNN backbone and a transformer architecture implementation.
@@
4.7 RTMDet
RTMDet [23] is a groundbreaking real-time object detection model designed to achieve optimal efficiency without compromising accuracy. It belongs to the family of fully convolutional single-stage detectors [90] such as YOLO series. RTMDet [23] operates as a one-stage detector, enabling it to swiftly recognize and localize objects in real-world scenarios such as autonomous driving, robotics, and drones. The model is engineered to push the boundaries of the YOLO series by introducing a new family of Real-Time Models for object Detection. Notably, RTMDet is capable of performing instance segmentation and rotated object detection, features that were previously unexplored.
Figure 8: One stage RTMDet architecture. Reproduced from [23].
RTMDet operates on a model architecture that emphasizes efficiency and compatibility in both the backbone and neck. The basic building block of the model consists of large-kernel depth-wise convolutions [91], which contribute to its ability to capture global context effectively. This architectural choice enhances the model’s capacity to recognize and localize objects with high precision. Furthermore, RTMDet incorporates soft labels during the calculation of matching costs in the dynamic label assignment process [92], leading to an improved accuracy. The combination of these architectural features and training techniques culminates in an object detector that achieves exceptional performance. RTMDet’s macro architecture follows the one-stage object detector paradigm, and it balances model depth, width, and resolution to optimize efficiency. Additionally, the model is designed to be versatile, allowing for easy extension to instance segmentation and rotated object detection tasks with minimal modifications.
In comparison to the state-of-the-art industrial detectors, the authors [23] demonstrate remarkable performance of RTMDet in terms of both speed and accuracy. The model achieves an impressive 52.8% AP on the COCO dataset while operating at over 300 frames per second (FPS) on an NVIDIA 3090 GPU. This outperforms the current mainstream industrial detectors, showcasing the superior parameter-accuracy trade-off of RTMDet. The model’s versatility is evident in its ability to deliver optimal performance across various application scenarios, offering different model sizes for different object recognition tasks.

### v4

@@ 4.2.1 You Only Look Once
The You Only Look Once (YOLO) framework [67] has emerged as a popular deep learning-based object detection algorithm that revolutionized real-time object detection tasks. It presents a unified approach to object detection by formulating it as a regression problem, allowing the model to predict bounding boxes and class probabilities of multiple objects in a single pass through the network.
In the early stages, YOLOv2 [78] introduced batch normalization and high-resolution classifiers to improve detection performance. YOLOv3 [43] further refined the architecture by incorporating skip connections and multi-scale prediction, enhancing both accuracy and localization capabilities. The introduction of CSPDarknet, SPP, PAN, and Mish activation in YOLOv4 [45] led to significant improvements in the network architecture. YOLOv5 [79], based on scaled-YOLOv4 [80], introduced anchor-free object detection and a new architecture. This version expanded the options available by providing models of varying sizes, allowing users to balance speed and accuracy according to their specific requirements (Figure 2).
@@
Some of the most known stable versions are YOLOv8 [29] and YOLO-NAS [82], which introduced the concept of neural architecture search (NAS) [83] to automatically design network architectures and achieve state-of-the-art performance in object detection tasks. More recently, YOLOv10 [28] introduced significant innovations to improve both efficiency and accuracy in real-time object detection. A key change is the move away from traditional non-maximum suppression (NMS) in favor of a consistent dual-assignment strategy. This change enables a more efficient end-to-end deployment with lower inference latency. Furthermore, YOLOv10 [28] incorporates Transformer-based enhancements, allowing for more effective attention mechanisms. Such a holistic model is designed to optimize various components for both computational efficiency and detection performance, resulting in reduced redundancy. This optimization leads to a model with 46% less latency, and 25% fewer parameters than YOLOv8 [29] while still achieving state-of-the-art results. The continuous evolution of the YOLO framework highlights the trade-offs between speed and accuracy, necessitating consideration of the specific application requirements when selecting an appropriate YOLO model.
4.2.2 Single-Shot Detection
@@
The Single-Shot Detector (SSD) [84] is an efficient and accurate object detection algorithm that introduces a unified framework for single-pass detection. In theory, SSD addresses the challenge of detecting objects at various scales including small targets [85] and aspect ratios, using predefined anchor boxes [86]. The SSD architecture consists of three main components: a base network, convolutional feature maps, and convolutional predictors (Figure 3). The base network, usually a pre-trained CNN [87], acts as a feature extractor, generating high-level feature maps with different spatial resolutions. These feature maps are then processed by convolutional predictors, which predict the presence and location of objects for the anchor boxes. Each predictor corresponds to a specific feature map and produces class scores and bounding box offsets. The anchor boxes, distributed densely throughout the feature map, serve as reference boxes for object detection. To capture objects at multiple scales, SSD employs feature maps from different stages of the base network. Higher-resolution feature maps are effective at detecting small objects, while lower-resolution ones are suitable for larger objects. The predictions from each feature map are combined to generate final class predictions and refined bounding box coordinates. During training, SSD uses a multitask loss function [88] that optimizes the model considering both localization loss (Smooth L1 loss) and classification loss (softmax loss) [89]. The localization loss penalizes the discrepancy between predicted and ground truth bounding box coordinates, while the classification loss encourages accurate class predictions. The loss is computed for positive and negative samples, including hard negatives based on confidence scores.
@@
Although SSD offers a good balance between accuracy and efficiency with its variable feature map resolutions and fixed anchor boxes, it does have limitations [90]. One limitation is the dependence on predefined anchor boxes, which may not adequately cover objects with extreme aspect ratios or unconventional shapes. Scaling the anchor boxes can help address this, but it introduces additional computational overhead. Another limitation is the challenge of handling tiny objects [91]. Since SSD relies on a limited number of feature maps, it may struggle to accurately detect small objects due to the loss of fine-grained details at higher resolutions. This can result in reduced localization accuracy and increased false negatives for small or densely packed objects. Furthermore, SSD’s fixed anchor boxes limit its ability to handle objects at arbitrary scales and aspect ratios.
4.2.3 RetinaNet Framework
RetinaNet is a one-stage object detection model that was introduced by Tsung-Yi Lin et al. [33] designed to address the extreme foreground–background class imbalance encountered during training of dense detectors [92]. In object detection, the goal is to detect objects of interest in an image and localize them by drawing bounding boxes around them. However, the vast majority of the image is typically background, and there are usually many more negative examples (background) than positive examples (objects of interest). This class imbalance can make it difficult for the detector to learn to distinguish between objects and background, and can lead to poor performance. RetinaNet is a single unified network composed of a backbone network and two task-specific subnetworks (Figure 4). The backbone is responsible for computing a convolutional feature map over an entire input image and is an off-the-shelf convolutional network. The first subnetwork is a dense prediction subnet that produces a fixed number of object detections of different scales and aspect ratios at each spatial location in a feature map. The second subnetwork is a set of class-specific subnets that further refine the predictions of the first subnet.
Figure 4: One-stage RetinaNet architecture. Reproduced from [21]. (a) ResNet [93] backbone. (b) Generation of multi-scale convolutional pyramid. This is attached to two subnetworks: (c) anchor box classification and (d) anchor box regression to ground-truth bounding box.
This model uses a novel focal loss function that down-weights the contribution of easy examples during training to focus on hard examples and prevents the vast number of easy background examples from overwhelming the detector. The focal loss is a dynamically scaled cross-entropy loss, where the scaling factor decays to zero as confidence in the correct class increases. This loss function is designed to address the mechanisms used by other detectors to address class imbalance, such as biased minibatch sampling and object proposal mechanisms, in a one-stage detection system directly via the loss function.
The obtained results showed good performance compared to previous one-stage and two-stage detectors, including the fastest reported R-CNN system, on the COCO dataset. It achieves state-of-the-art performance on both COCO detection tasks, with a better COCO test-dev average precision while running at 5 fps. This is a significant improvement over previous state-of-the-art techniques for training one-stage detectors, such as training with sampling heuristics or hard example mining. RetinaNet is also designed to be efficient and scalable. It uses an efficient in-network feature pyramid that allows it to detect objects at multiple scales and resolutions, and it uses anchor boxes to improve localization accuracy. The anchor boxes are predefined boxes of different scales and aspect ratios that are placed at each spatial location in the feature map. The dense prediction subnet predicts the offsets and scales of the anchor boxes to generate object detections.
4.2.4 CenterNet Framework
CenterNet [22] is another real-time object detection algorithm designed to operate in real time, with an average inference time of 270 ms using a 52-layer hourglass backbone and 340 ms using a 104-layer hourglass backbone per image according to the author (Figure 5). CenterNet is inspired by the architecture of CornerNet [32], which is based on a one-stage keypoint-based detector, while introducing several novel components and strategies to enhance its effectiveness.
Figure 5: One-stage CenterNet architecture. Reproduced from [22].
Unlike CornerNet, which detects object bounding boxes using pairs of keypoints, CenterNet introduces the concept of detecting each object as a triplet of keypoints. This innovation allows CenterNet to capture more comprehensive information about the objects, leading to improved detection performance while being fast for live inferencing. CenterNet incorporates two customized modules named cascade corner pooling and center pooling. These modules play crucial roles in enriching information collected by both top-left and bottom-right corners and providing more recognizable information at the central regions, respectively. These modules contribute to the improved performance of CenterNet by enhancing the detection of object keypoints and bounding boxes. In their benchmark study [22], the authors stated that this proposed architecture achieves significant improvements over existing one-stage detectors, typically an average precision of 47.0% on the MS-COCO dataset, outperforming the previously proposed one-stage detectors by at least 4.9%. Additionally, it demonstrates comparable performance to the two-stage detectors while maintaining a faster inference speed thanks to the effective reduction of incorrect bounding boxes, particularly for small objects. It achieves notable improvements in the detection of small objects, with an average precision (AP) improvements of 5.5% to 8.1% for different backbone configurations. This reduction in incorrect bounding boxes is attributed to the effectiveness of CenterNet in modeling center information using center keypoints.
4.2.5 RTMDet
RTMDet [30] is a groundbreaking real-time object detection model designed to achieve optimal efficiency without compromising accuracy. It belongs to the family of fully convolutional single-stage detectors [94] such as the YOLO series. RTMDet [30] operates as a one-stage detector, allowing it to quickly recognize and locate objects in real-world scenarios such as autonomous driving, robotics, and drones (Figure 6). The model is engineered to push the boundaries of the YOLO series by introducing a new family of real-time models for object detection. Notably, RTMDet is capable of performing instance segmentation and rotated object detection, features that were previously unexplored.
Figure 6: One-stage RTMDet architecture. Reproduced from [30].
RTMDet operates on a model architecture that emphasizes efficiency and compatibility in both the backbone and neck. The basic building block of the model consists of large-kernel depth-wise convolutions [95], which contribute to its ability to capture global context effectively. This architectural choice enhances the model’s capacity to recognize and localize objects with high precision. Furthermore, RTMDet incorporates soft labels during the calculation of matching costs in the dynamic label assignment process [96], leading to an improved accuracy. The combination of these architectural features and training techniques culminates in an object detector that achieves exceptional performance. RTMDet’s macro architecture follows the one-stage object detector paradigm, and balances the depth, width, and resolution of the model to optimize efficiency. Additionally, the model is designed to be versatile, allowing for easy extension to instance segmentation and rotated object detection tasks with minimal modifications.
In comparison to the state-of-the-art industrial detectors, the authors of [30] demonstrate the remarkable performance of RTMDet in terms of both speed and accuracy. The model achieves an impressive 52.8% AP on the COCO dataset while operating at over 300 frames per second (FPS) on an NVIDIA 3090 GPU. This outperforms the current mainstream industrial detectors, showcasing the superior parameter–accuracy trade-off of RTMDet. The model’s versatility is evident in its ability to deliver optimal performance across various application scenarios, offering different model sizes for different object recognition tasks.
4.3 Two-Stage Models
Region-Based CNNs
@@ 4.3 Two-Stage Models
The introduction of the Region-based Convolutional Neural Network (R-CNN) [97] marked a significant milestone in the development of object detection techniques [98], showcasing the substantial improvements that Convolutional Neural Networks (CNNs) can bring to detection performance. R-CNN introduced the concept of utilizing CNNs [99] in combination with a class-agnostic region proposal [100] module to transform object detection into a classification and localization problem. The detection process starts with a mean-subtracted input image, which is fed through the region proposal (RPN) module. This module employs techniques such as Selective Search [101] to identify regions within the image that have a higher likelihood of containing objects. Approximately 2000 object candidates are generated on the basis of this region proposal step. These candidates are then warped and passed through a CNN network, such as the widely used ImageNet [102], to extract a 4096-dimensional feature vector for each proposal. The feature vectors obtained from the CNN are then used as inputs for class-specific support vector machines (SVMs) [103], which have been trained beforehand. The SVMs generate confidence scores for each candidate region, aiding in the classification process. To refine the results, non-maximum suppression (NMS) is applied based on the intersection over union (IoU) and class information. Once the class has been identified, a trained bounding box regressor is employed to predict the precise bounding box coordinates, including the center coordinates, width, and height of the object (Figure 7).
@@
However, despite its groundbreaking contributions, R-CNN has several limitations [104, 105]. The R-CNN training process is complex and multistage. It involves pre-training the CNN with a large classification dataset, followed by fine-tuning on domain-specific images that undergo mean subtraction and warping to align with the proposals. The CNN classification layer is replaced with a randomly initialized N + 1-way classifier, where N represents the number of classes, and stochastic gradient descent (SGD) is utilized for optimization. Additionally, separate SVMs and bounding box regressors need to be trained for each class, adding to the computational complexity.
@@
Although R-CNN is capable of performing highly accurate object detection research, it suffered from slow inference times, taking approximately 47 s per image, and was resource-intensive in terms of both time and space [97]. Training R-CNN models on small datasets took days to complete, even with shared computations. These limitations sparked the need for further advancements in object detection algorithms that could address these challenges and improve overall efficiency.
4.4 Transformer-Based Architectures
4.4.1 DETR: End-to-End Transformer
@@
The advent of Transformer architectures has revolutionized the field of artificial intelligence by introducing a novel approach to processing sequential data. Unlike older neural network architectures, Transformers rely on self-attention mechanisms [58] to capture dependencies between input tokens, enabling them to effectively model long-range dependencies and capture complex patterns in the data. This ability to process sequences in parallel, rather than sequentially, has significantly improved the efficiency and effectiveness, especially of natural language processing (NLP) tasks [106]. The success of Transformers can be attributed to their ability to capture global context and relationships within the input data, making them particularly well suited for tasks that require an understanding of complex interdependencies.
@@
Later on, researchers translated the use of Transformers in NLP to computer vision, including object detection operation. One pioneering algorithm is DETR (DEtection TRansformer) [107], which belongs to the category of end-to-end object detection systems based on Transformers and bipartite matching loss for direct set prediction. Unlike traditional object detection methods, DETR does not fall into the conventional one-stage or two-stage detector categories. Instead, it introduces a novel approach by predicting all objects at once using a bipartite matching loss function, which uniquely assigns a prediction to a ground truth object and is invariant to a permutation of predicted objects. This unique approach simplifies the detection pipeline by eliminating the need for hand-designed components such as spatial anchors or non-maximal suppression [108], making it optimal for both accuracy and real-time processing. DETR achieves this by leveraging Transformers with parallel decoding, as opposed to autoregressive decoding with recurrent neural networks, which was the focus of previous work.
The architecture of DETR is simple yet highly effective. It comprises three main components: a CNN backbone for feature extraction [109], an encoder–decoder Transformer [106] for modeling relationships between feature representations of different detections, and a simple feed-forward network for making the final detection predictions (Figure 8). The process begins with the extraction of a lower-resolution activation map from the input image using a conventional CNN backbone. This activation map is then passed through the Transformer encoder, along with spatial positional encodings that are added to queries and keys at every multi-head self-attention layer. The decoder receives queries, output positional encodings (object queries), and encoder memory, and produces the final set of predicted class labels and bounding boxes through multiple multi-head self-attention and decoder–encoder attention (FFNs). The simplicity and modularity of the DETR architecture make it easily implementable in any deep learning framework that provides a common CNN backbone and a Transformer architecture implementation.
@@
4.4.2 Grounding DINO
Grounding DINO [112] is a novel object detection model that merges the DINO (DEtection with INterpolation Optimization) architecture with grounded pre-training, aiming for effective open-set object detection. This integration allows the model to not only perform classical object detection but to identify arbitrary objects guided by textual descriptions or category names, expanding its utility in scenarios where predefined class categories are insufficient. The Grounding DINO model relies on a Transformer-based architecture [58] that introduces several innovations for fusing visual and language data. The core components (summary in Figure 9) include the following: (1) Visual Backbone through the integration of convolutional backbones to extract visual features from the input image, similarly to standard object detection models. These features are then embedded into the Transformer architecture. (2) Language Integration is a critical feature of Grounding DINO incorporating natural language descriptions for object identification similarly to open-world object detection [113]. To achieve this, it uses a pre-trained language model to encode textual inputs, such as category names or phrases. The encoded language vectors are then tightly integrated with visual features through a multi-step fusion process. (3) Query Selection and Cross-Modality Decoder: The model employs a feature enhancer that refines visual features using language input to guide the search for relevant objects. A language-guided query selection mechanism enables the identification of regions of interest in the image by selecting queries that match linguistic descriptions. Finally, a cross-modality decoder integrates these selected queries, refining detection and enabling accurate bounding box predictions and classifications even for previously unseen categories.
Figure 9: Grounding DINO architecture. Reproduced from [112].
Grounding DINO’s evaluation showcases its capabilities in both conventional and zero-shot detection tasks. It achieves 52.5 average precision (AP) on the COCO zero-shot benchmark [26], indicating its strong ability to generalize without training data from the dataset. The model sets a new standard on the ODinW zero-shot benchmark [114], obtaining a mean 26.1 AP for open-world detection scenarios where the model must handle unseen object categories. Grounding DINO also demonstrates strong performance on referring expression comprehension, meaning it can accurately detect objects when prompted by attribute-based descriptions, as seen in tasks like RefCOCO and RefCOCOg [115].